OpenStudy (anonymous):

please help me guys out there.. i'll compare ur answer to mine. Find the solution set using Quadratic Formula ax2 +bx+c=0 Solve for the unknown. (t-4)(T is not equal to 5)(T-6)=6 ()

jhonyy9 (jhonyy9):

if consider t-4=0 --- t=4 if t-6=6 --- t=6+6=12